OpenStudy (anonymous):

In a reaction,5g ethyl acetate is hydrolysed per litre in the presence of dil. HCl in 300 min. If the reaction is of first order, and the initial concentration of ethyl acetate is 22g/L, then calculate the rate constant of the reaction.

OpenStudy (chmvijay):

yes put your old equation and solve it where are u facing the problem?

OpenStudy (anonymous):

@chmvijay Mass is given not the concentration that is the problem

OpenStudy (chmvijay):

concenteration is what amount of substance per liter of solvent right here also the same that is nothing but g/L =fyou just find it :)

OpenStudy (anonymous):

@chmvijay i will try out wait

OpenStudy (anonymous):

@chmvijay I am geting the answer as 4.93*10^-3 whereas the answer is supposed to be 8.6*10^-4

OpenStudy (chmvijay):

can you show ur work up

OpenStudy (anonymous):

\[k=\frac{ 2.303 }{ 300 }\log_{10} \frac{ 22 }{ 5 } = 4.9395\times10 ^{-3}\]

OpenStudy (anonymous):

@chmvijay

OpenStudy (anonymous):

i did it with a calculator.... @Frostbite

OpenStudy (chmvijay):

22 and 5 are not here the initial and final concentration right

OpenStudy (chmvijay):

yes i got it

OpenStudy (chmvijay):

it should be k=2.303/300 log 22/17 since initial concentration is 22g/L when it hydrolysis in 300 minutes mean the final concentration will be 22-5 =17g/Litre now you do it you get ur answer :)

OpenStudy (anonymous):

@chmvijay genius!! Thanks a lot......are u an iitian?

OpenStudy (chmvijay):

I am Proud to be an INDIAN :)

OpenStudy (anonymous):

m asking about iit.....

OpenStudy (chmvijay):

NOOOO :(

OpenStudy (anonymous):

ok anyways thanks!!

OpenStudy (chmvijay):

Ur Welcome :)

OpenStudy (chmvijay):

Dont think only IITians are GREAT and is the only hallmark source of knowledge:) take it positively

OpenStudy (anonymous):

i know i was just curious @chmvijay do not get me wrong actually i am an aspirant so i was asking........

OpenStudy (chmvijay):

hey its OK, no problem , my words will help you in future :)thank you

OpenStudy (dean.shyy):

This may help: http://is.gd/vCfBYe